OpenStudy (anonymous):

Solve the matrix equation: [-7,-3 2,1] X=[2,6 6,9]

OpenStudy (anonymous):

\[X=\left[\begin{matrix}x_1 & x_2 \\ x_3 & x_4\end{matrix}\right]\] using matrix multiplication \[-7x_1-3x_3=2\]\[2x_1+x_3=6\]\[-7x_2-3x_4=6\]\[2x_2+x_4=9\] two systems of 2 equations with 2 knowns, solve for x1,x2,x3,x4

OpenStudy (anonymous):

im not sure what to do next..can u explain more

OpenStudy (anonymous):

not much more to explain you should be able to do systems of equations if you're learning this

OpenStudy (anonymous):

are you learning about inverses of matrices?